Jul 02 DAY shift summary

Richard Kessler (rkessler@fnal.gov)
Fri, 2 Jul 1999 15:38:20 -0500 (CDT)

Shift Summary for DAY shift, July 02, 1999

On Shift:RK EJR

Hours of beam delivered at all intensities:
Hours of beam delivered at nominal intensity:
Hours of E832 data (special runs not included):
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: 15 minutes to change tapes

========================================================================
Runs ended this shift:
Run 13791 - 7.3M events - BEAM_832 (10.8K vac00) ==> tapes full ends run
Run 13792 - 5.0M events - BEAM_832 ( 7.4K vac00) ==> ktev1 fire ends run

Tapes brought to FCC this shift: QKK73-75 and B025

========================================================================
Accesses during shift: we resisted all temptations to make an access.

========================================================================
Overview of shift:
Dimples Donuts appear at the beginning of shift, and they
are loaded with sugar, fat and cholesterol. One of shift
crew indulges; the other does not.

Smooth E832 data taking.

During run 13792 we got a few ktevana errors; looking at the
event dump shows unknown crate 17 showing up between the two
REG ADC crates ?? This corresponds to KQ 4X, but the KQ4X
DYC is now completely bypassed, dart-cables and all. So there
is still something generating a phony crate 17 in stream 6.

And now the excitement begins. We smelled smoke in the control room
shortly after noon and eventually found smoke billowing out the back
of KTEV1. Special thanks to the well calibrated nostrils of AB and EJR.
KTEV1 was turned off immediately and this ends run 13792 with no EOF
marks. Tapes pulled out at 15% full. Tevatron goes down at about the
same time.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
7:30 Munster tripped off causing 50% humidity in hall.
L.Nelson drops by to turn it back on.

13792 124 12:25 Ends with ktev1 smoking and turned off.
No EOF marks

13:00 SGI is contacted about ktev1.
13:42 SGI calls back and says 1/2 hour to arrive.
14:50 SGI guys arrive.

13794 14:00 use BEAM_832_NOKTEV1 configuration.
Note that BAF1 and DDD1 vetos were unplugged.
ktev1 also removed from targets.lis file.
Waiting for beam when run starts.

48 15:17 Beam arrives at 5E12. DA running smoothly.
15:30 SGI guy returns to PREP for more parts.

========================================================================
Follow up Notes on Previous Shift

ETOT and HCC are fine (after lots of swapping last shift).

========================================================================
Items to Follow up on Next Shift: ktev1 ??

- must fill current tapes [QKK079-80 and QKB027]
using BEAM_832_NOKTEV1:
- DO NOT USE ANY OTHER CONFIG WITHOUT CONTACTING EXPERTS.

- DO NOT DO ONSPLIT BACKUP WIHTOUT CONTACTING EXPERTS ...
ktev1 may not be up so lots of data could be missing.